OpenStudy (anonymous):

Can someone help with this question please. Find the probability. What is the probability that a card drawn from a deck of 52 cards is not a 10? A. 12/13 B. 9/10 C. 1/13 D. 1/10

Directrix (directrix):

There are 4 tens in the deck. There are 52 - 4 = 48 cards in the deck which are not tens. P(not a "10" on one draw) = number of desired outcomes/number of possible outcomes. P(~10) = 48/52 = ?